# Taming Cross-Domain Representation Variance in Federated Prototype Learning with Heterogeneous Data Domains

### v1

@@ Abstract
Keywords Federated Prototype Learning  $\cdot$ Domain Heterogeneity
@@ 1 Introduction
Figure 1: Illustration of federated learning with heterogeneous data domains. The Vanilla column depicts the local feature distribution of the standard FedPL approach, obtaining average local and global prototypes directly. Proposed method showcased in the adjacent column yields a larger inter-class distance and a reduced intra-class distance. Note that without capturing variance information, even for hard domains, local averaged prototypes for each class can be well distinguished while the feature vectors are still mixed up. Both methods illustrate noticeable variations in domain characteristics across datasets, as detailed in section 4.
To address the aforementioned issue of data heterogeneity, numerous FL techniques have been devised [20, 5, 38, 29, 14, 12]. While these methods have indeed improved convergence, their focus on non-IID scenarios remains largely restricted. Typically, they assume that the data across clients pertain to a single domain, attributing non-IID distribution to label skew alone. Yet, in more realistic scenarios, clients gather data according to their unique preferences, making it impractical to assume identical domain origins for local data. Instead, the data often stems from heterogeneous domains, resulting in varied feature distributions. In contrast to the plethora of research on label skew-induced heterogeneity within the same domain, studies exploring domain heterogeneity are comparatively scarce. Existing literature [10] indicates that simplistic learning approaches applied across varied domain data yield limited generalizability, as global models fail to adequately capture the nuances of diverse domains. To overcome this challenge, employing prototypes—representative mean values of feature vectors with identical semantics that encapsulate class characteristics—may offer a potential solution in FL settings plagued by domain heterogeneity.
@@
Considering the challenges posed by unequal learning due to domain diversity, we introduce a novel method, termed FedPLVM (short for Federated Prototype Learning with Variance Mitigation). FedPLVM devises two main mechanisms. Firstly, we develop a dual-level prototype clustering mechanism that adeptly captures variance information, a significant improvement over previous methodologies that rely on averaging local training samples’ representations to derive local prototypes. Direct averaging inadequately represents the extensive sample variance inherent in these areas, resulting in unrecognized disparate learning difficulties. This discrepancy leads to proficient learning outcomes in simpler domains but decayed performance in more complex ones. Our local-level clustering generates multiple local clustered prototypes within each domain. To further mitigate increased communication costs and privacy concerns arising from transferring a comprehensive set of diverse prototypes for each class from every client, we employ global-level prototype clustering on the server side. This strategy reduces the number of prototypes each client requires. Secondly, by capturing the variance information through clustered prototypes, we design an innovative $\alpha$-sparsity prototype loss to enhance the training process. Unlike existing methods that utilize a single prototype for each class, our approach, which employs clustered prototypes, increases the risk of overlapping feature representations across different classes in challenging clients. To prevent the intermingling of feature distributions from ‘hard’ clients, instead of simply maximizing the feature-level distance between each local instance and the prototypes of other classes, we refine this distance by elevating it to an $\alpha$ power, with $\alpha$ being a value between 0 and 1. This modification effectively repels other prototypes, thereby introducing greater sparsity in the inter-class feature distributions. Moreover, we incorporate an intra-class similarity correction term to lessen the feature-level distance among intra-class samples, thereby concentrating the $\alpha$-sparsity prototype loss on amplifying the feature-level distance across inter-class samples. Collectively, these two strategies empower FedPLVM to reasonably leverage variance information, thus equilibrating fairness between ‘hard’ and ‘easy’ learning domains and enhancing overall learning performance. Consequently, FedPLVM emerges as a reliable approach for Federated Learning in contexts characterized by heterogeneous data domains. Our main contributions are outlined as follows:
@@ 2 Related Work
FL aims to train a global model through collaboration among multiple clients while preserving their data privacy. FedAvg [21], the pioneering work in FL, demonstrates the advantages of this approach in terms of privacy and communication efficiency by aggregating local model parameters to train a global model. A significant challenge in FL is data heterogeneity, often manifested as non-IID (independently and identically distributed) data. Subsequent research, following FedAvg, has primarily focused on addressing data heterogeneity to enhance training performance in FL environments. Specifically, studies such as [18, 2, 30] have improved performance by incorporating a global penalty term to mitigate discrepancies. Other works, e.g., [32, 23], have sought to maximize feature-level agreement between local and global models to further boost performance. Recent works [10, 32, 19, 37] have explored data heterogeneity arising from client-specific domain diversity, while overlooking the challenges of unequal learning across different domains. In this paper, we propose a novel approach using dual-level prototype clustering to capture essential local variance information. Additionally, we introduce a new $\alpha$-sparsity loss, specifically designed to tackle the challenges of learning in diverse domains, thereby facilitating the development of a more generalizable global model.
2.2 Prototype Learning.
@@
Prototype learning has been extensively explored in various tasks, such as transfer learning [27, 13], few-shot learning [31], zero-shot learning [11], and unsupervised learning [34]. The concept of a prototype in this context refers to the average feature vectors of samples within the same class. In the FL literature, prototypes serve to abstract knowledge while preserving privacy. Specifically, approaches like FedProc [23] and FedProto [32] focus on achieving feature-wise alignment with global prototypes. FedPCL [33] employs prototypes to capture knowledge across clients, constructing client representations in a prototype-wise contrastive manner using a set of pre-trained models. FPL [10] highlights the use of cluster-based and unbiased global prototypes to tackle the challenges in FL where clients possess domain-diverse data. Our study addresses a similar issue as FPL but with distinct emphases. While FPL concentrates on the disparities in the number of clients across domains, aiming to mitigate the bias in the overall model caused by domains with more clients, our work focuses on the intrinsic learning challenges that vary across domains.
2.3 Contrastive Learning.
@@
Contrastive learning is a promising self-supervised learning technique. The work by [25] constructs pairs of positives and negatives for each sample and applies the InfoNCE loss to compare these pairs. Another work by [15] extends contrastive learning from self-supervised to fully supervised settings, utilizing both label information and contrastive methods. Additionally, some studies [16, 36] integrate contrastive learning into local training to enhance performance in FL. Instead of the conventional InfoNCE loss, our approach introduces a new $\alpha$-sparsity loss, aiming to further reduce similarity among inter-class sample features while amplifying similarity among intra-class samples.
@@ 3 Preliminary
3.1 Regular FL Scenario.
Consider the classic FL scenario: there exist $K$ clients and one server, which aims to assist all clients to train a common machine learning model without sharing their private data, denoted by $\mathcal{D}_{k}=\{\mathbf{x}_{i},y_{i}\}^{N_{k}}_{i}$ for client $k$. Formally, the global objective of FL can be formulated as:
@@
3.2 Domain Shift in FL.
In the simplest FL setting, the label and feature distribution are the same between two clients $i$ and $j$, formally $P_{i}(y)=P_{j}(y)$ and $P_{i}(x|y)=P_{j}(x|y)$. Extending such a scenario to the heterogeneous data setting brings us to the point of domain shift, also denoted as feature non-IID setting. Domain shift is caused by distinctive feature distributions among clients, that is $P_{i}(x|y)\neq P_{j}(x|y)$, though their label space is still the same.
3.3 Federated Prototype Learning.
To handle the FL domain shift problem, previous works [32, 10] divide the classification network into two parts: feature extractor and classifier. The feature extractor $h:\mathbb{R}^{V}\rightarrow\mathbb{R}^{D}$ maps a sample $\mathbf{x}\in\mathbb{R}^{V}$ into the feature space and generates the corresponding feature vector $\mathbf{z}=h(\mathbf{x})\in\mathbb{R}^{D}$. Then the classifier $f:\mathbb{R}^{D}\rightarrow\mathbb{R}^{M}$ outputs the $M$-class prediction $f(\mathbf{z})=f(h(\mathbf{x}))\in\mathbb{R}^{M}$ given the feature vector $\mathbf{z}$. The intuition of FedPL is adjusting the feature extractor to generate a consistent feature distribution among different domains. Similar to the global shareable model, the straightforward solution computes the average feature vector of all samples belonging to the same class for further processing, called the prototype.
Due to the distributed property of FL, we cannot collect all samples via the server, which makes the above-mentioned idea a two-step procedure:
@@ 5 Experiments
Implement Details. We employ the ResNet10 [9] as our backbone model, configuring the feature vectors’ dimension to 512. The optimization is done using the SGD optimizer, employing a learning rate of $0.01$, momentum of $0.5$, and a weight decay of $51e-$ 5. For Digit-5, Office-10 and DomainNet, we use 5, 4 and 6 clients respectively. The client data is independent and identically distributed (i.i.d.) and non i.i.d. results can be found in Supplementary Material. It is important to note that each client operates within distinct data domains, meaning different datasets. For Digit-5 and Office-10, each client possessed 100 training samples and 1000 test samples. Global communication rounds are fixed at $T=50$ for Digit-5 and $T=80$ for Office-10. Each local training epoch consists of $E=2$ iterations. We maintain default hyper-parameter values: $\tau=0.07$, $\alpha=0.25$, and $\lambda=100$. As for DomainNet, we followed the setup in FedPCL using a 10-class subset. Each client employs 300 training samples and all test samples (approximately 1000, vary on domains). $\lambda=1$ and $T=200$ for DomainNet. The batch size is set at 32 for all datasets. Details regarding hyper-parameter settings will be elaborated in section 5.2. For fair comparisons, we conduct each setting for 5 experiments and report the average result.
Table 1: Test accuracy on Digit-5. Avg means average results among all clients. Details in section 5.1.
@@
Table 2: Test accuracy on Office-10. Details in section 5.1.
@@
Table 3: Test accuracy on DomainNet. Details in section 5.1.

### v2

@@ 1 Introduction
Figure 1: Illustration of federated learning with heterogeneous data domains. The Vanilla column depicts the local feature distribution of the standard FedPL approach, obtaining average local and global prototypes directly. Proposed method showcased in the adjacent column yields a larger inter-class distance and a reduced intra-class distance. Note that without capturing variance information, even for hard domains, local averaged prototypes for each class can be well distinguished while the feature vectors are still mixed up. Both methods illustrate noticeable variations in domain characteristics across datasets, as detailed in Fig. 4.
To address the aforementioned issue of data heterogeneity, numerous FL techniques have been devised [20, 5, 41, 29, 14, 12]. While these methods have indeed improved convergence, their focus on non-IID scenarios remains largely restricted. Typically, they assume that the data across clients pertain to a single domain, attributing non-IID distribution to label skew alone. Yet, in more realistic scenarios, clients gather data according to their unique preferences, making it impractical to assume identical domain origins for local data. Instead, the data often stems from heterogeneous domains, resulting in varied feature distributions.
@@
Considering the challenges posed by unequal learning due to domain diversity, we introduce a novel method, termed FedPLVM (short for Federated Prototype Learning with Variance Mitigation). FedPLVM devises two main mechanisms. Firstly, we develop a dual-level prototype clustering mechanism that adeptly captures variance information, a significant improvement over previous methodologies that rely on averaging local training samples’ representations to derive local prototypes. Our local-level clustering generates multiple local clustered prototypes within each domain. To further mitigate increased communication costs and privacy concerns arising from transferring a comprehensive set of diverse prototypes for each class from every client, we employ global-level prototype clustering on the server side. Secondly, by capturing the variance information through clustered prototypes, we design an innovative $\alpha$-sparsity prototype loss to enhance the training process. To prevent the intermingling of feature distributions from ‘hard’ clients, instead of simply maximizing the feature-level distance between each local instance and the prototypes of other classes, we refine this distance by elevating it to an $\alpha$ power, with $\alpha$ being a value between 0 and 1. This modification effectively repels other prototypes, thereby introducing greater sparsity in the inter-class feature distributions. Moreover, we incorporate an intra-class similarity correction term to lessen the feature-level distance among intra-class samples, thereby concentrating the $\alpha$-sparsity prototype loss on amplifying the feature-level distance across inter-class samples. Collectively, these two strategies empower FedPLVM to reasonably leverage variance information, thus equilibrating fairness between ‘hard’ and ‘easy’ learning domains and enhancing overall learning performance. Consequently, FedPLVM emerges as a reliable approach for Federated Learning in contexts characterized by heterogeneous data domains. Our main contributions are outlined as follows:
@@ 2 Related Work
Federated Learning. FL aims to train a global model through collaboration among multiple clients while preserving their data privacy. FedAvg [21], the pioneering work in FL, demonstrates the advantages of this approach in terms of privacy and communication efficiency by aggregating local model parameters to train a global model. A significant challenge in FL is data heterogeneity, often manifested as non-IID (independently and identically distributed) data. Subsequent research, following FedAvg, has primarily focused on addressing data heterogeneity to enhance training performance in FL environments. Specifically, studies such as [18, 2, 30] have improved performance by incorporating a global penalty term to mitigate discrepancies. Other works, e.g., [32, 23], have sought to maximize feature-level agreement between local and global models to further boost performance. Recent works [10, 32, 19, 40] have explored data heterogeneity arising from client-specific domain diversity, while overlooking the challenges of unequal learning across different domains. In this paper, we propose a novel approach using dual-level prototype clustering to capture essential local variance information. Additionally, we introduce a new $\alpha$-sparsity loss, specifically designed to tackle the challenges of learning in diverse domains, thereby facilitating the development of a more generalizable global model.
@@
Prototype Learning. Prototype learning has been extensively explored in various tasks, such as transfer learning [27, 13], few-shot learning [31], zero-shot learning [11], and unsupervised learning [34]. The concept of a prototype in this context refers to the average feature vectors of samples within the same class. In the FL literature, prototypes serve to abstract knowledge while preserving privacy. Specifically, approaches like FedProc [23] and FedProto [32] focus on achieving feature-wise alignment with global prototypes. FedPCL [33] employs prototypes to capture knowledge across clients, constructing client representations in a prototype-wise contrastive manner using a set of pre-trained models. FPL [10] highlights the use of cluster-based and unbiased global prototypes to tackle the challenges in FL where clients possess domain-diverse data. Our study addresses a similar issue as FPL but with distinct emphases. While FPL concentrates on the disparities in the number of clients across domains, aiming to mitigate the bias in the overall model caused by domains with more clients, our work focuses on the intrinsic learning challenges that vary across domains.
@@
Contrastive Learning. Contrastive learning is a promising self-supervised learning technique. The work by [25] constructs pairs of positives and negatives for each sample and applies the InfoNCE loss to compare these pairs. Another work by [15] extends contrastive learning from self-supervised to fully supervised settings, utilizing both label information and contrastive methods. Additionally, some studies [16, 39] integrate contrastive learning into local training to enhance performance in FL. Instead of the conventional InfoNCE loss, our approach introduces a new $\alpha$-sparsity loss, aiming to further reduce similarity among inter-class sample features while amplifying similarity among intra-class samples.
@@ 3 Preliminary
Regular FL Scenario. Consider the classic FL scenario: there exist $K$ clients and one server, which aims to assist all clients to train a common machine learning model without sharing their private data, denoted by $\mathcal{D}_{k}=\{\mathbf{x}_{i},y_{i}\}^{N_{k}}_{i}$ for client $k$. Formally, the global objective of FL can be formulated as:
@@
Domain Shift in FL. In the simplest FL setting, the label and feature distribution are the same between two clients $i$ and $j$, formally $P_{i}(y)=P_{j}(y)$ and $P_{i}(x|y)=P_{j}(x|y)$. Extending such a scenario to the heterogeneous data setting brings us to the point of domain shift, also denoted as feature non-IID setting. Domain shift is caused by distinctive feature distributions among clients, that is $P_{i}(x|y)\neq P_{j}(x|y)$, though their label space is still the same.
Federated Prototype Learning. To handle the FL domain shift problem, previous works [32, 10] divide the classification network into two parts: feature extractor and classifier. The feature extractor $h:\mathbb{R}^{V}\rightarrow\mathbb{R}^{D}$ maps a sample $\mathbf{x}\in\mathbb{R}^{V}$ into the feature space and generates the corresponding feature vector $\mathbf{z}=h(\mathbf{x})\in\mathbb{R}^{D}$. Then the classifier $f:\mathbb{R}^{D}\rightarrow\mathbb{R}^{M}$ outputs the $M$-class prediction $f(\mathbf{z})=f(h(\mathbf{x}))\in\mathbb{R}^{M}$ given the feature vector $\mathbf{z}$. The intuition of FedPL is adjusting the feature extractor to generate a consistent feature distribution among different domains. Similar to the global shareable model, the straightforward solution computes the average feature vector of all samples belonging to the same class for further processing, called the prototype. Due to the distributed property of FL, we cannot collect all samples via the server, which makes the above-mentioned idea a two-step procedure:
@@ 5 Experiments
Implement Details. We employ the ResNet10 [9] as our backbone model, configuring the feature vectors’ dimension to 512. The optimization is done using the SGD optimizer, employing a learning rate of $0.01$, momentum of $0.5$, and a weight decay of $51e-$ 5. For Digit-5, Office-10 and DomainNet, we use 5, 4 and 6 clients respectively. The client data is independent and identically distributed (i.i.d.) and non i.i.d. results can be found in Supplementary Material. It is important to note that each client operates within distinct data domains, meaning different datasets. For Digit-5 and Office-10, each client possessed 100 training samples and 1000 test samples. Global communication rounds are fixed at $T=50$ for Digit-5 and $T=80$ for Office-10. Each local training epoch consists of $E=2$ iterations. We maintain default hyper-parameter values: $\tau=0.07$, $\alpha=0.25$, and $\lambda=100$. As for DomainNet, we followed the setup in FedPCL using a 10-class subset. Each client employs 300 training samples and all test samples (approximately 1000, vary on domains). $\lambda=1$ and $T=200$ for DomainNet. The batch size is set at 32 for all datasets. Details regarding hyper-parameter settings will be elaborated in Sec. 5.2. For fair comparisons, we conduct each setting for 5 experiments and report the average result.
Table 1: Test accuracy on Digit-5. Avg means average results among all clients. Details in Sec. 5.1.
@@
Table 2: Test accuracy on Office-10. Details in Sec. 5.1.
@@
5.1 Performance Comparison
We compare our proposed method with the state-of-the-art methods using the Digit-5, Office-10 and DomainNet datasets, as detailed in Tab. 1, Tab. 2 and Tab. 6. Our method demonstrates significant improvements in average accuracy over baseline methods for both datasets. A closer examination of domain-specific results reveals a more pronounced enhancement in performance on domains that are more challenging to learn. For instance, within the Digit-5 dataset, our method achieves a $5.3\%$ increase in accuracy for the SVHN domain, which is more difficult, and a $0.58\%$ increase for the MNIST dataset, considered easier. This aligns with the goal of our method, which is to address the disparate learning challenges across various domains, enhancing fairness and aiding in performance improvement in harder domains. Similar improvements can also be observed in Office-10 and DomainNet in Sec. C of Appendix.
5.2 Ablation Study
5.2.1 Impact of Dual-Level Prototype Generation.
In this subsection, we focus on examining the impact of dual-level clustered prototypes, as demonstrated by the results in Tab. 3. The first row illustrates outcomes derived from calculating local prototypes by averaging features of samples within the same class, and global prototypes through the direct averaging of these local prototypes. This approach, utilizing straightforward averaging for both levels, yields the least effective performance. This is attributed to its failure to capture variance information, a non-trivial aspect in this context. Moreover, adopting a solely global clustering approach does not significantly enhance performance. This is attributed to the fact that while global clusters incorporate inter-domain variance, they overlook the critical aspect of high intra-domain sample variance, particularly in domains that are challenging to learn. To elucidate this, we present a t-SNE visualization analysis in Fig. 3 comparing the three prototype generation methods. Our approach fosters a more generalizable decision boundary, as illustrated in the visualization. This ability to effectively capture variance at both local and global levels is key to why our dual-level clustering method outperforms the others.
Table 3: Comparison on prototype generation methods. Variance means the average distance from the normalized feature vector of one sample to its corresponding class feature center (i.e. the averaged prototype). Results are then used for visualization in Fig. 3. Details in Sec. 5.2.1.
Figure 3: Visualization of different prototype generation methods. The first row averages feature vectors locally and averages local prototypes globally. The second row averages feature vectors locally and clusters local prototypes globally. The last row (ours) clusters feature vectors locally and clusters local clustered prototypes globally. The last column Total is the visualization of mixing the feature vectors from all datasets. Details in Sec. 5.2.1.
To further explore alternative methods, we examine the efficacy of directly transferring all local prototypes to each client without employing any aggregation or clustering techniques. This approach preserves variance information but raises significant privacy concerns, unlike our dual-level clustering method, which offers enhanced privacy protection. As demonstrated in Tab. 4, directly transferring all local prototypes yields performance comparable to our dual-cluster approach. However, it requires transmitting approximately five times as many prototypes in each training round.
Table 4: Comparison between w/o and w/ global clustering. w/o means the server distributes all local clustered prototypes to the clients for local training. Avg ##\## of prototypes is the average number of prototypes each client receives from the server during each global round. Details in Sec. 5.2.1.
5.2.2 Impact of $\alpha$-Sparsity Prototype Loss.
Table 5: Comparison on components of $\alpha$-sparsity prototype loss. Contrast and Correction stand for the contrastive and corrective loss term in the total $\alpha$-sparsity loss respectively. Avg is the average accuracy result for all clients. Details in Sec. 5.2.2.
To evaluate the specific impact of our proposed $\alpha$-sparsity prototype loss, we conduct experiments comparing both contrastive and corrective loss terms. Results in Tab. 5 demonstrates that employing the contrast term led to a $1.46\%$ improvement in final average accuracy, while the correction term resulted in a $0.99\%$ enhancement. Combining all components yields the best performance, showcasing a $2.29\%$ improvement. Our investigation also focuses on the sparsity parameter $\alpha$, depicted in Fig. 4. We find that varying $\alpha$ within the range of $(0,1)$ consistently outperforms the baseline setting with $\alpha=1$. After careful consideration of numerical stability, we opt for an $\alpha$ value of $0.250$.
Figure 4: Impact of $\alpha$ sparsity and $\lambda$ prototype loss weight. The left figure shows the accuracy of two selected datasets and the average accuracy among all clients with different $\alpha$. The right figure shows the effects of different $\lambda$ for both FPL and our proposed approach. Details in Sec. 5.2.2.
5.3 Impact of Temperature $\tau$.
To assess the impact of the contrastive temperature ($\tau$) on our model’s performance, we conduct the following experiment. The results, depicted in Fig. 5, demonstrate that our method consistently surpasses the baselines across a range of temperatures (refer to Tab. 1 and Tab. 2 for details, where FPL exhibits the highest baseline performance at $66.88\%$ and $58.55\%$ respectively). Further analysis indicates that an optimal temperature setting for our model on Digit-5 is $\tau=0.070$. Another results on Office-10 identifies an optimal temperature setting for our model at $\tau=0.045$, while we maintain $\tau=0.070$ for stability in performance comparisons. It is noteworthy that temperatures either significantly higher or lower than this value lead to training difficulties due to numerical instability.
Figure 5: Impact of $\tau$. The left figure shows the impact on the Digital-5 dataset while the right figure shows on the Office-10 dataset. Average means the average test accuracy with variance among all clients. Details in Sec. 5.3.
6 Conclusion
Acknowledgements
This work is supported in part by NSF under grants 2033681, 2006630, 2044991, 2319780.
Appendix A Supplementary Material Overview
The supplementary material is organized into the following sections:
B: Detailed algorithm of our proposed method, FedPLVM.
C: Further results of our experiments.
D: Additional quantitative results comparing FedPLVM and other baseline methods under the label non-i.i.d. setting.
E: Additional quantitative results comparing different clustering algorithms for the prototype clustering procedure of FedPLVM.
F: Additional quantitative results comparing FedPLVM and FPL[10] under the unbalanced client domain distribution setting.
G: Additional quantitative results employing differential privacy technology in conjunction with FedPLVM.
H: Additional quantitative results comparing different global prototype generation methodologies of FedPLVM.
I: Additional quantitative results analyzing the effectiveness of local prototypes clustering on different domains.
J: Discussion on the wider range of scenario of our method, including limitation and broader impact.
Appendix B Detailed Algorithm of FedPLVM
We summarize our proposed method, FedPLVM in the Alg. 1.
Algorithm 1 FedPLVM
1:Input: Communication rounds $T$, local training epochs $E$, number of classes $M$, number of clients $K$, private dataset $\mathcal{D}_{k}=\{\mathbf{x}_{i},y_{i}\}^{N_{k}}_{i}$
2:Output: Global model $w^{T+1}$
3:Server Aggregation:
4:for $t=1,2,...,T$ do
5:     for $k=1,2,...,K$ do
6:         Collect local models and clustered prototypes $w^{t}_{k,E+1},\{\mathcal{P}^{m}_{k}\}_{m=1}^{M}\leftarrow$ Local Update $(k,w^{t},\mathcal{G})$
7:     end for
8:     Aggregate collected prototypes $\{\mathcal{P}^{m}\}_{m=1}^{M}$
9:     Generate global clustered prototypes $\mathcal{G}$ in Eq. 4
10:     Aggregate global model $w^{t+1}=\sum_{k=1}^{K}\frac{N_{k}}{N}w^{t}_{k,E+1}$
11:end for
12:Local Update$(k,w^{t},\mathcal{G})$:
1:$w^{t}_{k,1}\leftarrow w^{t}$
2:for $e=1,2,...,E$ do
3:     Update $w^{t}_{k,e+1}$ from $w^{t}_{k,e}$ using $\mathcal{G}$ by Eq. 10
4:end for
5:Compute local feature vectors $\{h(\mathbf{x}_{i})\}_{i=1}^{N_{k}}$ for $\mathcal{D}_{k}$
6:Generate local clustered prototypes $\{\mathcal{P}^{m}_{k}\}_{m=1}^{M}$ in Eq. 3
7:Return $w^{t}_{k,E+1},\{\mathcal{P}^{m}_{k}\}_{m=1}^{M}$
Appendix C Further results on DomainNet
Due to page limitation, we put some further experimental results in Sec. 5.1 here:
Table 6: Test accuracy on DomainNet. Details in Sec. 5.1.